# Multi-Agent Reinforcement Learning with Communication-Constrained Priors

### v1

@@ 1 Introduction
In multi-agent reinforcement learning (MARL) with partial observations, collaboration poses a significant challenge [oroojlooy2023review]. Communication is one of the effective measures to improve the learning of cooperative policy in MARL [zhu2024survey], widely applied in scenarios such as autonomous driving [chen2024communication, xu2024distributed] and cooperative drones [chafii2023emergent, kim2024cooperative]. However, real-world scenarios are far from ideal and communication between agents often faces various constraints, specifically: (1) limited communication bandwidth, meaning only a limited amount of message can be transmitted, and (2) lossy communication, where transmitted message may be subject to interference, delay, loss, and other issues.
Much of the research on communication-constrained multi-agent reinforcement learning (MARL) focuses on the limited bandwidth issue [wang2019learning, lee2021learning, hu2024learning, zhou2024semantic]. In this setting, it is assumed that the communication channel is ideal, which means that transmission is real-time and lossless. Therefore, these methods typically only need to focus on how to effectively allocate communication resources (e.g., communication bandwidth, communication medium) to promote cooperation among agents. For example, compressing the communication information to extract and transmit parts beneficial to cooperation [liu2020multi], and systematically allocating communication mediums to avoid competition, among other problems [kim2019learning].
However, in most real-world scenarios, communication links are uncertain, and lossy communication is more prevalent. In this setting, existing work typically addresses two types of issue: noise interference and communication delay. Approaches to addressing noise interference typically involve modeling the unknown noise distribution and constructing learnable processes to adaptively optimize cooperative policy [freed2020communication]. Another more common issue is communication delay, which frequently occurs in wireless network environments, which refers to the non-real-time transmission of message, It has been shown to impact the performance of multi-agent behaviors [zhang2020succinct, yuan2023dacom]. Approaches to addressing communication delay focus primarily on how to remove the impact of delayed message on cooperative policy, such as constructing communication buffer [zhang2020succinct] and determining when to communicate [yuan2023dacom]. Nevertheless, the above methods are all based on ideal assumptions about communication delay, and may not be applicable to more complex and unknown real-world scenarios such as underwater and caves. The reasons are as follows: (1) These methods lack scalability due to the lack of consideration of the common characteristics of lossy communication in different unknown scenarios. (2) These methods lack robustness due to the lack of consideration of the dilemma of promoting the relevance of effective communication messages and suppressing the relevance of lossy communication information.
@@ 2.1 MARL with Communication Bandwidth Constraints
Research on communication bandwidth constraints has primarily focused on how to achieve efficient interaction among multiple agents under limited bandwidth resources. zhang2019efficient proposed the VBC method, which reduces communication overhead by constraining the variance of information exchanged between agents during the training phase, thereby eliminating noise from the messages. kim2019learning addressed the issue of medium contention in the information transmission process, proposing the SchedNet framework. By incorporating a MAC protocol from the wireless communication field, this framework alleviates bandwidth pressure and addresses the scheduling problem of constrained agents. mao2020learning introduced the Gated-ACML algorithm, which uses a gating mechanism to filter messages. It adaptively prunes unnecessary information by determining whether a message is beneficial using a threshold, thus reducing the number of messages exchanged. hu2023event proposed the ETCNet framework, which converts limited bandwidth into a penalty threshold for event-triggered strategies, improving communication efficiency in multi-agent systems by sending messages only when necessary.
These methods essentially perform effective compression on messages, which aligns closely with the information bottleneck [tishby2015deep] principle—eliminating task-irrelevant redundant information and extracting task-relevant refined information. Such compression approaches have been validated to exhibit a certain degree of generalization capability in similar tasks [yang2025learning]. However, compression is not equivalent to robustness [yann2020learning, wang2022pac]. In environments with unstable or lossy communication, these compressed methods may not be effective. For instance, unstable communication is highly prone to losing critical messages, leading to a significant degradation in the quality of policies.
@@ 2.2 MARL with Lossy Communication
Research on lossy communication has primarily focused on the robustness of multi-agent systems under non-ideal communication conditions, such as environmental noise, transmission delays, and packet loss. freed2020communication proposed a novel differentiable communication method using a randomized message encoding scheme, where it mathematically equates discrete communication channels to simulated channels with additive noise, enabling gradient backpropagation through these channels. kim2019message proposed the Message-Dropout training method, based on the concept of Dropout, which randomly drops communication messages from other agents during training to improve robustness against communication errors during execution. zhang2020succinct introduced the TMC method, which leverages temporal locality to reduce redundant messages and incorporates a message buffering mechanism to enhance robustness against packet loss, making it suitable for bandwidth-constrained and packet-loss-prone network environments. yuan2023dacom proposed DACOM, an adaptive delay-aware multi-agent communication model in which agents can learn to schedule waiting times for messages from other agents, which is suitable for delay-sensitive tasks and high-latency scenarios.
@@ 3 Problem Formulation
Considering a fully collaborative multi-agent task where each agent is in a partially observable and communicative environment, it can be modeled as a decentralized partially observable Markov decision process (Dec-POMDP) [albrecht2024multi] with communication, which is defined by the following 9-tuple, . is denoted by the set of agents. is denoted by the set of environmental states. is observation set of all agents, and is an observation set for agent , which is determined by observation function . is the set of agents’ joint action space, where is denoted by the action space of agent . is the global reward signal shared by the agents. is denoted by the transition probability. is the discount factor.
@@ 4.1 Communication-Constrained Priors Modeling
where is the parameter determined by the environment, and is the part of the state most relevant to agents and . For different environments, can be defined manually or obtained through pre-training of binary classification tasks. In the context of specific learning, there are several approaches: () When addressing specific and stable communication-constrained environments, one can estimate the priors of communication links through sampling or empirical data, enabling policy learning to adapt as closely as possible to these specific conditions. () In the face of diverse and non-stable communication-constrained environments, designing more diverse priors for communication links can allow policy learning to cover multiple exceptional scenarios. Specifically, message-dropout [kim2019message] is considered a case of this prior modeling, where messages are randomly masked with a certain probability to adapt to communication constraints. Therefore, incorporating such priors helps distinguish more effectively between lossy and lossless messages.
@@ 4.2.1 MI between Messages and Behaviors
MI is often used to improve multi-agent collaboration in MARL [wang2019learning, li2022pmic]. It is a theory that measures the correlation between different variables. When extended to communication-constrained MARL, a basic requirement is to measure the correlation between messages and agent behaviors. So we have
@@ 4.2.2 Du-MIE for Constrained Communication
Note that the basic definitions of communication-constrained prior and MI, we thus formulate a natural objective propose for communication-constrained MARL: maximize the MI between lossless messages and agent behaviors (to enhance their positive correlations) while minimizing the MI between lossy messages and agent behaviors (to mitigate their negative correlations). However, exact computation of MI is highly challenging, as it necessitates simultaneously determining both the joint probability distribution and the marginal distributions of the variables involved [belghazi2018mutual]. Inspired by the Dual Mutual Information Estimatior (Du-MIE) [li2022pmic], we construct a Du-MIE for constrained communication to estimate the above objective. The Jensen-Shannon MI estimator based on Jensen-Shannon divergence (JSD) [hjelm2019learning] estimates the lower bound of for maximization and the Contrastive Log-ratio Upper Bound (CLUB) [cheng2020club] estimates the upper bound of for minimization, based on replay buffer with lossless messages and lossy messages respectively. Specifically, JSD can be defined as
where is message-action joint distribution for agent , is the product of the marginals. is the softplus function. is a discriminator function modeled by a neural network with parameters . and are obtained by joint sampling from the replay buffer with lossless messages , while is sampled independently from the same buffer. By minimizing the loss , we can make closely approximate the lower bound of [hjelm2019learning]. On the contrary, CLUB can be defined as
@@ 4.3 Communication-Constrained MARL
In order to characterize the impact of the upper and lower bounds of mutual information on multi-agent policy learning, inspired by reward shaping [li2022pmic], we reshape the global reward as follows:
@@
This reward shaping can be combined with different MARL algorithms. Taking CTDE-based MARL algorithms as an example, the results of the combination are presented in Algorithm 1. Both policy-based [lowe2017multi] and value-based [rashid2020monotonic] methods can derive a global temporal difference (TD) loss, which is presented as follows:
@@ 5.1 Experimental Setup
In the experimental setup, we integrate the proposed algorithm framework with MADDPG [lowe2017multi] to form Communication-Constrained MADDPG (CC-MADDPG) as the primary validation target. It is then compared with four baselines: MAIC [yuan2022multi], Full-Communication MADDPG (FC-MADDPG), Dropout-MADDPG, and the standard MADDPG, operating without inter-agent communication. We adopt the Multi-Agent Particle Environments (MPEs) [lowe2017multi] as benchmarks. To simulate communication constraints, we employ the following two distinct models,
Markov-Based Communication (MBC): The Markov model [zhang2020succinct] assumes that the state of a system at any time is determined by a state transition probability matrix, which includes one noiseless state and multiple lossy states—the more the number of lossy states, the higher the loss probability. In this experiment, the transition probability matrices are set with dimensions of and , corresponding to loss levels of light, medium, and heavy, respectively.
Distance-Based Communication (DBC): This approach [yuan2023dacom] simulates signal attenuation in real-world environments (e.g., underwater or cave-like) based on inter-agent distances. Specifically, it sets a distance threshold to determine the degree of communication constraint, where smaller distances lead to higher message loss rates. In this experiment, the distance threshold are set with and , corresponding to loss levels of light, medium and heavy, respectively.
@@ 5.2.2 Impact of Communication Constraint Priors
Light DBC | -128.720.3 | -119.029.4 | |||
| Medium DBC | -127.619.2 | -98.030.5 | ||||
| Heavy DBC | -128.020.6 | -107.033.3 | ||||
Light DBC | 75.576.6 | 84.284.1 | |||
| Medium DBC | 73.874.5 | 107.5109.1 | ||||
| Heavy DBC | 62.170.1 | 80.286.7 | ||||
@@ 6 Conclusion
This work is supported in part by National Natural Science Foundation of China (62192783, 62276128, 62506157, 62276142), Jiangsu Natural Science Foundation (BK20243051), Jiangsu Science and Technology Major Project (BG2024031), the Fundamental Research Funds for the Central Universities (14380128), Nanjing International/Hong Kong, Macao and Taiwan Science and Technology Cooperation Plan (202308031) and the Collaborative Innovation Center of Novel Software Technology and Industrialization.

### v2

@@ 1 Introduction
In multi-agent reinforcement learning (MARL) with partial observations, collaboration poses a significant challenge [19]. Communication is one of the effective measures to improve the learning of cooperative policy in MARL [31], widely applied in scenarios such as autonomous driving [4, 24] and cooperative drones [3, 12]. However, real-world scenarios are far from ideal and communication between agents often faces various constraints, specifically: (1) limited communication bandwidth, meaning only a limited amount of message can be transmitted, and (2) lossy communication, where transmitted message may be subject to interference, delay, loss, and other issues.
Much of the research on communication-constrained multi-agent reinforcement learning (MARL) focuses on the limited bandwidth issue [22, 14, 10, 30]. In this setting, it is assumed that the communication channel is ideal, which means that transmission is real-time and lossless. Therefore, these methods typically only need to focus on how to effectively allocate communication resources (e.g., communication bandwidth, communication medium) to promote cooperation among agents. For example, compressing the communication information to extract and transmit parts beneficial to cooperation [16], and systematically allocating communication mediums to avoid competition, among other problems [11].
However, in most real-world scenarios, communication links are uncertain, and lossy communication is more prevalent. In this setting, existing work typically addresses two types of issue: noise interference and communication delay. Approaches to addressing noise interference typically involve modeling the unknown noise distribution and constructing learnable processes to adaptively optimize cooperative policy [7]. Another more common issue is communication delay, which frequently occurs in wireless network environments, which refers to the non-real-time transmission of message, It has been shown to impact the performance of multi-agent behaviors [29, 27]. Approaches to addressing communication delay focus primarily on how to remove the impact of delayed message on cooperative policy, such as constructing communication buffer [29] and determining when to communicate [27]. Nevertheless, the above methods are all based on ideal assumptions about communication delay, and may not be applicable to more complex and unknown real-world scenarios such as underwater and caves. The reasons are as follows: (1) These methods lack scalability due to the lack of consideration of the common characteristics of lossy communication in different unknown scenarios. (2) These methods lack robustness due to the lack of consideration of the dilemma of promoting the relevance of effective communication messages and suppressing the relevance of lossy communication information.
@@ 2.1 MARL with Communication Bandwidth Constraints
Research on communication bandwidth constraints has primarily focused on how to achieve efficient interaction among multiple agents under limited bandwidth resources. Zhang et al. [28] proposed the VBC method, which reduces communication overhead by constraining the variance of information exchanged between agents during the training phase, thereby eliminating noise from the messages. Kim et al. [11] addressed the issue of medium contention in the information transmission process, proposing the SchedNet framework. By incorporating a MAC protocol from the wireless communication field, this framework alleviates bandwidth pressure and addresses the scheduling problem of constrained agents. Mao et al. [18] introduced the Gated-ACML algorithm, which uses a gating mechanism to filter messages. It adaptively prunes unnecessary information by determining whether a message is beneficial using a threshold, thus reducing the number of messages exchanged. Hu et al. [9] proposed the ETCNet framework, which converts limited bandwidth into a penalty threshold for event-triggered strategies, improving communication efficiency in multi-agent systems by sending messages only when necessary.
These methods essentially perform effective compression on messages, which aligns closely with the information bottleneck [21] principle—eliminating task-irrelevant redundant information and extracting task-relevant refined information. Such compression approaches have been validated to exhibit a certain degree of generalization capability in similar tasks [25]. However, compression is not equivalent to robustness [6, 23]. In environments with unstable or lossy communication, these compressed methods may not be effective. For instance, unstable communication is highly prone to losing critical messages, leading to a significant degradation in the quality of policies.
@@ 2.2 MARL with Lossy Communication
Research on lossy communication has primarily focused on the robustness of multi-agent systems under non-ideal communication conditions, such as environmental noise, transmission delays, and packet loss. Freed et al. [7] proposed a novel differentiable communication method using a randomized message encoding scheme, where it mathematically equates discrete communication channels to simulated channels with additive noise, enabling gradient backpropagation through these channels. Kim et al. [13] proposed the Message-Dropout training method, based on the concept of Dropout, which randomly drops communication messages from other agents during training to improve robustness against communication errors during execution. Zhang et al. [29] introduced the TMC method, which leverages temporal locality to reduce redundant messages and incorporates a message buffering mechanism to enhance robustness against packet loss, making it suitable for bandwidth-constrained and packet-loss-prone network environments. Yuan et al. [27] proposed DACOM, an adaptive delay-aware multi-agent communication model in which agents can learn to schedule waiting times for messages from other agents, which is suitable for delay-sensitive tasks and high-latency scenarios.
@@ 3 Problem Formulation
Considering a fully collaborative multi-agent task where each agent is in a partially observable and communicative environment, it can be modeled as a decentralized partially observable Markov decision process (Dec-POMDP) [1] with communication, which is defined by the following 9-tuple, . is denoted by the set of agents. is denoted by the set of environmental states. is observation set of all agents, and is an observation set for agent , which is determined by observation function . is the set of agents’ joint action space, where is denoted by the action space of agent . is the global reward signal shared by the agents. is denoted by the transition probability. is the discount factor.
@@ 4.1 Communication-Constrained Priors Modeling
where is the parameter determined by the environment, and is the part of the state most relevant to agents and . For different environments, can be defined manually or obtained through pre-training of binary classification tasks. In the context of specific learning, there are several approaches: () When addressing specific and stable communication-constrained environments, one can estimate the priors of communication links through sampling or empirical data, enabling policy learning to adapt as closely as possible to these specific conditions. () In the face of diverse and non-stable communication-constrained environments, designing more diverse priors for communication links can allow policy learning to cover multiple exceptional scenarios. Specifically, message-dropout [13] is considered a case of this prior modeling, where messages are randomly masked with a certain probability to adapt to communication constraints. Therefore, incorporating such priors helps distinguish more effectively between lossy and lossless messages.
@@ 4.2.1 MI between Messages and Behaviors
MI is often used to improve multi-agent collaboration in MARL [22, 15]. It is a theory that measures the correlation between different variables. When extended to communication-constrained MARL, a basic requirement is to measure the correlation between messages and agent behaviors. So we have
@@ 4.2.2 Du-MIE for Constrained Communication
Note that the basic definitions of communication-constrained prior and MI, we thus formulate a natural objective propose for communication-constrained MARL: maximize the MI between lossless messages and agent behaviors (to enhance their positive correlations) while minimizing the MI between lossy messages and agent behaviors (to mitigate their negative correlations). However, exact computation of MI is highly challenging, as it necessitates simultaneously determining both the joint probability distribution and the marginal distributions of the variables involved [2]. Inspired by the Dual Mutual Information Estimatior (Du-MIE) [15], we construct a Du-MIE for constrained communication to estimate the above objective. The Jensen-Shannon MI estimator based on Jensen-Shannon divergence (JSD) [8] estimates the lower bound of for maximization and the Contrastive Log-ratio Upper Bound (CLUB) [5] estimates the upper bound of for minimization, based on replay buffer with lossless messages and lossy messages respectively. Specifically, JSD can be defined as
where is message-action joint distribution for agent , is the product of the marginals. is the softplus function. is a discriminator function modeled by a neural network with parameters . and are obtained by joint sampling from the replay buffer with lossless messages , while is sampled independently from the same buffer. By minimizing the loss , we can make closely approximate the lower bound of [8]. On the contrary, CLUB can be defined as
@@ 4.3 Communication-Constrained MARL
In order to characterize the impact of the upper and lower bounds of mutual information on multi-agent policy learning, inspired by reward shaping [15], we reshape the global reward as follows:
@@
This reward shaping can be combined with different MARL algorithms. Taking CTDE-based MARL algorithms as an example, the results of the combination are presented in Algorithm 1. Both policy-based [17] and value-based [20] methods can derive a global temporal difference (TD) loss, which is presented as follows:
@@ 5.1 Experimental Setup
In the experimental setup, we integrate the proposed algorithm framework with MADDPG [17] to form Communication-Constrained MADDPG (CC-MADDPG) as the primary validation target. It is then compared with four baselines: MAIC [26], Full-Communication MADDPG (FC-MADDPG), Dropout-MADDPG, and the standard MADDPG, operating without inter-agent communication. We adopt the Multi-Agent Particle Environments (MPEs) [17] as benchmarks. To simulate communication constraints, we employ the following two distinct models,
Markov-Based Communication (MBC): The Markov model [29] assumes that the state of a system at any time is determined by a state transition probability matrix, which includes one noiseless state and multiple lossy states—the more the number of lossy states, the higher the loss probability. In this experiment, the transition probability matrices are set with dimensions of and , corresponding to loss levels of light, medium, and heavy, respectively.
Distance-Based Communication (DBC): This approach [27] simulates signal attenuation in real-world environments (e.g., underwater or cave-like) based on inter-agent distances. Specifically, it sets a distance threshold to determine the degree of communication constraint, where smaller distances lead to higher message loss rates. In this experiment, the distance threshold are set with and , corresponding to loss levels of light, medium and heavy, respectively.
@@ 5.2.2 Impact of Communication Constraint Priors
| Task Scenario | Testing Environment | Priors Type | |
| dropout-0.2 | Test-Matched | ||
| Simple_Spread | Light DBC | -128.720.3 | -119.029.4 |
| Medium DBC | -127.619.2 | -98.030.5 | |
| Heavy DBC | -128.020.6 | -107.033.3 | |
| Simple_Reference | Light DBC | 75.576.6 | 84.284.1 |
| Medium DBC | 73.874.5 | 107.5109.1 | |
| Heavy DBC | 62.170.1 | 80.286.7 | |
@@ 6 Conclusion
This work is supported in part by National Natural Science Foundation of China (62192783, 62276128, 62506157, 62276142), Jiangsu Natural Science Foundation (BK20243051), Jiangsu Science and Technology Major Project (BG2024031), the Fundamental Research Funds for the Central Universities (14380128, KG202514), Nanjing International/Hong Kong, Macao and Taiwan Science and Technology Cooperation Plan (202308031) and the Collaborative Innovation Center of Novel Software Technology and Industrialization.